find the 54th term of an A.P 100,97,94---
Answer by Theo(13342)

100, 97, 94
d = -3

formula for AP is:

An = A1 + (n-1)*d

A1 = 100
d = -3
n = 54

formula becomes:

A54 = 100 + (53*-3) which becomes:

A54 = 100 - 159 which becomes:

A54 = -59